A discount store sold plastic cups for $3.25 each and ceramic cups for $4.50 each. If 500 cups were sold for a total of $2090 how many cups of each type was sold? B. What was the dollar value of each type of cup sold?
Answer by rfer(12731)

p+c=500 c=500-p
3.25p+4.50c=2090
3.25p+4.50(500-p)=2090
3.25p+2250-4.50p=2090
-1.25p=-160
p=128 plastic cups
c=500-128=372 ceramic cups
------------------
P=128*3.25=$416
c=372*4.50=$1674